Determine the energy of a photon with a wavelength of 357 nm. (h = 6.626 × 10⁻³⁴ J • s and c = 3.00 × 10⁸ m/s)
Answer #1

Step 1: Explanation

We know, E= hv where h = planck's constant = 6.626 × 10-34 J-s

frequency (v) = C/λ

and C = 3 × 108 m/s =speed of light

so the equation becomes

or, E= hv

or, E= hc / λ

Step 2: Calculation

Given, wavelength = 357 nm = 357 × 10-9 m [ because 1 nm = 10-9 m ]

So by using above equation

E= hc / λ

=> E= ( 6.626 × 10-34 J-s × 3 × 108 m/s )   / 357 × 10-9 m = 5.57  ×10-19 J
Answer #2

6846513513132